### v1

# Curriculum Negative Mining For Temporal Networks

Abstract
Temporal networks are effective in capturing the evolving interactions of networks over time, such as social networks and e-commerce networks. In recent years, researchers have primarily concentrated on developing specific model architectures for Temporal Graph Neural Networks (TGNNs) in order to improve the representation quality of temporal nodes and edges. However, limited attention has been given to the quality of negative samples during the training of TGNNs. When compared with static networks, temporal networks present two specific challenges for negative sampling: positive sparsity and positive shift. Positive sparsity refers to the presence of a single positive sample amidst numerous negative samples at each timestamp, while positive shift relates to the variations in positive samples across different timestamps. To robustly address these challenges in training TGNNs, we introduce Curriculum Negative Mining (CurNM), a model-aware curriculum learning framework that adaptively adjusts the difficulty of negative samples. Within this framework, we first establish a dynamically updated negative pool that balances random, historical, and hard negatives to address the challenges posed by positive sparsity. Secondly, we implement a temporal-aware negative selection module that focuses on learning from the disentangled factors of recently active edges, thus accurately capturing shifting preferences. Extensive experiments on 12 datasets and 3 TGNNs demonstrate that our method outperforms baseline methods by a significant margin. Additionally, thorough ablation studies and parameter sensitivity experiments verify the usefulness and robustness of our approach. Our code is available at https://github.com/zziyue83/CurNM.
keywords:
Temporal Graph Neural Networks , Negative sampling , Curriculum learning , Disentanglement learning
journal: Neural Networks
[…]
[1]organization=Department of Economics, University of California, Berkeley,city=Berkeley, state=California, country=United States
[…]
[2]organization=Big Graph Center, School of Computer and Computing Science, Hangzhou City University,city=Hangzhou, state=Zhejiang, country=China
[…]
[3]organization=State Key Laboratory of Blockchain and Data Security, Zhejiang University,city=Hangzhou, state=Zhejiang, country=China
[…]
[4]organization=Hangzhou High-Tech Zone (Binjiang) Institute of Blockchain and Data Security,city=Hangzhou, state=Zhejiang, country=China
1 Introduction
Temporal networks have gained popularity in recent years due to their capacity to capture the evolving nature inherent in diverse domains, such as social networks [18, 21, 28], recommendation systems [1, 16], financial transactions [22], transport [26], and political networks [6, 11, 19]. Unlike their static counterparts, these networks support additions, deletions, and changes in the features of nodes and edges over time by modeling a sequence of time-stamped events. To best leverage this temporal information, Temporal Graph Neural Networks (TGNNs) [8, 16, 24, 25, 27, 30, 33] have been developed.
Similar to static graph models, TGNNs [24, 33] employ negative sampling techniques to prevent the models from being overwhelmed by a disproportionally high number of negative over positive samples and enhance the models’ efficiency. The most straightforward approaches often adopted are random sampling and recent sampling [24, 33]. However, temporal networks introduce two unique challenges that particularly undermine these approaches: positive sparsity and positive shift. Positive sparsity emerges as the inclusion of a temporal dimension significantly increases the negative-to-positive ratio by the number of timestamps, compared to static graphs. Consequently, random sampling in this context tends to produce a multitude of irrelevant and easily distinguishable negative samples, offering little to no value. Positive shift, on the other hand, occurs as the relationships between nodes change over time; for instance, in recommendation systems, users’ preferences are likely to evolve [16]. Ignoring this factor may result in training models on outdated information, particularly in datasets spanning long periods.
Nonetheless, the negative sampling problem in temporal network learning has been seldom explored. While such efforts are limited for TGNNs, there is a wealth of strategies developed for static graphs. Whether by adopting alternative distributions [4, 32, 36] or employing model-aware methods to select negatives that closely resemble positives  [5, 12, 17, 38, 39], these existing strategies seek to enhance models’ performance by focusing on more informative negative samples, specifically those that closely resemble positive ones, during training. However, applying these strategies directly to temporal networks is problematic due to positive sparsity. In particular, given the scarcity of positive samples, using negative samples that are too similar to positives can skew the classification boundary excessively towards the negatives. This can lead to confusion within the model, particularly during early epochs, and result in poor performance. Thus, a more gradual transition from easy to hard negatives is needed. Moreover, positive shift induces a negative shift, rendering existing methods for static graphs ineffective for temporal networks. This highlights the need for components that can incorporate recent information and update dynamically over time. Till now, ENS [7] is the only negative sampling method specifically designed for temporal networks. While targeting temporal networks, ENS adopts a heuristic, global negative mining strategy instead of an adaptive, personalized approach, leading to unstable performance. Additionally, it suffers from low efficiency due to sampling from all historical edges.
To address these challenges, we introduce the CurNM method. Overall, CurNM is a curriculum learning strategy that dynamically adjusts the learning difficulty of negative samples based on the model’s performance, thereby addressing the prevalent issue of steep difficulty curves in existing methods. Subsequently, we incorporate two components to address the unique characteristics of temporal networks. Firstly, to address positive sparsity, we construct a negative pool for selection, which includes a mix of random samples and historical neighbors, with hard negatives added if the model shows strong learning progress. Starting with a relatively small negative pool also helps maintain computational efficiency. Secondly, we design a negative selection function that leverages temporal-aware disentanglement to capture positive shifts and individual node preferences, which further enhances the model’s granularity.
Our key contributions are summarized as follows:
  • 1.
    We identify two crucial challenges of negative sampling in temporal networks, i.e., positive sparsity and positive shift.
  • 2.
    An efficient curriculum learning framework on negative sampling is tailored for temporal networks to handle these two challenges.
  • 3.
    Elaborate experiments on 12 temporal networks and 3 TGNNs show that our CurNM can consistently boost the link prediction performance compared to baselines.
  • 4.
    CurNM is highly portable and can be integrated as a plug-in into any TGNN with minimal modifications.
2 Related Work
2.1 Temporal Graph Learning
Motivated by the significant progress of deep learning in Computer Vision [15] and Natural Language Processing [20], learning dense embeddings for nodes has become a de facto standard paradigm in graph learning [9, 23], unifying different graph tasks into a graph representation learning task. Subsequently, Graph Neural Networks [10, 14, 31], which take both graph structure and node features into account, replaced skip-gram models [9, 23] for graph representations. Influenced by the advances of static graph representation learning, TGNNs have been proposed to handle the evolving dynamics of temporal networks [29], updating node embeddings as time elapses.
Existing TGNNs can be classified into discrete-time and continuous-time TGNNs based on their temporal natures [13]. On the one hand, discrete-time TGNNs [8, 22, 25, 27] focus on the sequential dynamics of a series of graph snapshots, where the timespan of a snapshot depends largely on the graph domain. The consequent discrete-time TGNNs include a combination of Recurrent Neural Networks and Graph Neural Networks. DySAT [25], a typical discrete-time TGNN, first learns the structural representations of nodes within each graph snapshot and then aggregates across multiple snapshots using a temporal self-attention layer.
On the other hand, continuous-time TGNNs focus on the sequential interactions of nodes, thereby offering more fine-grained temporal information. Various design paradigms exist for continuous-time TGNNs since temporal networks can be seen as sequential dynamics of either nodes or a graph. JODIE [16] utilizes two mutually updated RNNs to update user and item embeddings, alongside a temporal attention layer that projects user embeddings after a certain time since their last interaction. Conversely, TGN [24] presents a generic framework on continuous temporal networks, simulating both RNNs [16, 30] and TGNNs [33] with its flexible memory mechanism and GNNs. Recent state-of-the-art continuous-time TGNNs [34, 37] have further explored the self-attention mechanism on temporal networks, inspired by the advancement of Transformer networks [2]. However, TGNNs typically employ random and recent negative sampling [24, 33], with little attention given to the specific study of negative sampling strategies within this domain.
2.2 Hard Negative Sampling
Negative sampling strategies [4, 5, 12, 17, 32, 36, 38, 39] have been abundant in graph representation learning such as collaborative filtering and knowledge graphs. These strategies are traditionally categorized into two types: static and model-aware negative sampling.
Figure 1: An illustration of our CurNM method for temporal networks, which progressively increases the difficulty of sampled negatives. $(a)$ illustrates the pipeline of TGNNs in temporal network link prediction, where our CurNM adds a Negative Pool and a Negative Selection module to the basic random sampling approach. $(b)$ illustrates the construction of a fixed-size, dynamically updated negative pool. $(c)$ shows the selection of negatives from this pool based on the candidates’ disentangled distances and their temporal proximity to the interactions. $(d)$ depicts the final negative samples used for training TGNNs, which include selected negatives from $(c)$ combined with random negatives. $(e)$ illustrates the process of selecting negatives in random sampling, where irrelevant negatives are likely to be chosen at any stage of training.
Static negative sampling strategies [20] select negative samples based on a pre-defined probability distribution. The simplest and most common strategy is random sampling, which uniformly selects negative samples from uninteracted nodes. However, random sampling falls short in ensuring the quality of the samples. In response, alternative distributions have been proposed. For instance, NNCF [4] and NCEPLRec [32] adopt item-popularity-based sampling distributions to preferentially select popular items as negative samples, thereby mitigating the widespread popularity bias in recommender systems [3]. Despite these advancements, all these methods fail to adequately account for distribution variations across different networks, thus leading to unsatisfactory performance.
To tackle this challenge, model-aware negative sampling strategies leverage models’ specific information to oversample negatives that closely resemble positive samples, thereby tightening models’ decision boundaries and enhancing accuracy. DNS [36] pioneers this approach by dynamically choosing negative samples with the highest prediction scores. NSCaching [38] builds on this by keeping a cache of hard negatives, whereas SRNS [5] introduces a variance-based criterion to reduce the risk of false negatives. Moreover, MixGCF [12] synthesizes hard negatives by infusing positive information into negative samples. DENS [17] splits embeddings into relevant and irrelevant factors and selects negatives that only differ in relevant factors from the positives. ANS [39] combines factor-aware and augmentation techniques from these strategies; it separates the hard and easy factors of negative items and synthesizes new negatives by augmenting the easy factors. While these aforementioned strategies have shown encouraging outcomes, their direct application to TGNNs is disappointing due to their oversight of temporal networks’ unique characteristics, particularly positive shift and positive sparsity.
To the best of our knowledge, ENS [7] is the only negative sampling method specifically designed for temporal networks. ENS selects negative samples by aiming to meet a difficulty target composed of the proportion of historical neighbors and average temporal proximity. While ENS tries to tackle positive sparsity by increasing the difficulty target by a predefined amount each epoch, its heuristic approach fails to adapt to the varying needs of different models and datasets. Moreover, although it addresses positive shift by considering temporal proximity, this method requires evaluating all historical edges to ensure accuracy, thereby consuming substantial resources. Finally, ENS does not account for individual node preferences; it indiscriminately prioritizes all nodes that have historically appeared, without distinguishing historical neighbors, and fails to recognize that interactions between users and items are often driven by certain key factors.
3 Problem Statement
Temporal Network
This section formulates the problem of negative sampling in temporal networks. Let $G(V,E,T,X)$ be a temporal network where $V$ and $E$ denote the set of nodes and edges, $T$ denotes the timestamps of interactions, and $X$ denotes the node features. We denote the set of interactions by $O^{+}=\{(u,v^{+},t)|u,v^{+}\in V\text{ and }t\in T\}$ where each tuple $(u,v^{+},t)$ records an interaction between two nodes $u,v^{+}$ that occurs at time $t$. By convention, we call $u$ source node and $v^{+}$ destination node. These pairs form positive training samples. Conversely, we define the complement of $O^{+}$ as $O^{-}=\{(u,v^{-},t)|u,v^{-}\in V,t\in T\text{ and }(u,v^{-},t)\notin O^{+}\}$. This set contains all node pairs $u,v^{-}$ that are not connected by an edge at time $t$ and is used to generate negative samples. Our goal is to define a negative sampling strategy $f$ that selects negative samples $(u,v^{-},t)$ from $O^{-}$ to enhance the performance of existing TGNNs.
Historical Neighbor
$u,v^{\prime}$ are called neighbors at $t$ if $(u,v^{\prime},t)\in O^{+}$. If there exists $(u,v^{\prime\prime},t^{\prime})\in O^{+}$ such that $t^{\prime}\leq t$, $v^{\prime\prime}$ is called a historical neighbor of $u$ at $t$. In our context, we further restrict historical neighbors to nodes that are not neighbors of $u$ at the current time, i.e., $(u,v^{\prime\prime},t)\notin O^{+}$. For each $(u,t)\in(V,T)$, we maintain a set $H_{u,t}=\{(v_{1},t_{1}),(v_{2},t_{2}),...,(v_{n},t_{n})\}$ that records all unique historical neighbors of $u$ at $t$ and their respective interaction timestamps. In cases where $v_{i}$ interacts with $u$ multiple times before $t$, we use the latest timestamp for $t_{i}$. Historical neighbors in our method are then drawn from this set $H_{u,t}$.
4 Methodology
Overall, CurNM is a curriculum learning strategy that progressively increases the difficulty of selected negatives based on performance on the validation set, effectively addressing the steep difficulty curves found in existing methods. Within this framework, we further develop two components. Firstly, to tackle positive sparsity, we establish a negative pool consisting of historical neighbors and random negatives, with hard negatives included when the model exhibits sustained improvements. Secondly, we use a temporal-aware disentangling technique to score the information of candidate negatives in the negative pool, emphasizing those that differ from positives only in the most critical dimensions, thus capturing positive shifts and individual preferences. These informative negative samples are then fetched for training TGNNs to ensure their performance and robustness.
This framework is illustrated in Figure 1, and the detailed workflow is presented in Algorithm 1.
Algorithm 1 The training process with CurNM.
0:  Training set $O^{+}=\{(u,v^{+},t)\}$, Negative set $O^{-}=\{(u,v^{-},t)\}$, Temporal network model $F$.
1:  for $e=1$ to $E_{0}$ do
2:     Get the embeddings of nodes by $F$.
3:     Sample a mini-batch $O^{+}_{\text{batch}}$ from $O^{+}$.
4:     Update the difficulty level $\pi_{e}$ by Eq. 1.
5:     for all $(u,v^{+},t)\in O^{+}_{\text{batch}}$ do
6:        Check $\mathcal{C}_{\pi_{e}}$ whether to cache training negatives by Eq. 5.
7:        if not $\mathcal{C}_{\pi_{e}}$ then
8:           Get the negative pool $V_{u}^{-}$ by Eq. 4.
9:        else
10:           Get the negative pool $V_{u}^{-}$ by Eq. 7.
11:        end if
12:        Disentangle the relevant factors by Eq. 8 and Eq. 9.
13:        Disentangle the irrelevant factors by Eq. 10.
14:        Get the scores of node pairs by Eq. 13.
15:        Select informative node pairs $O^{-}_{\text{sel}}$ by Eq. 2.
16:        Update embeddings by Eq. 3.
17:        if $\mathcal{C}_{\pi_{e}}$ then
18:           Update $V_{u}^{hard}$ by Eq. 6.
19:        end if
20:     end for
21:  end for
4.1 Curriculum Learning
To address the learning difficulty caused by positive sparsity, the curriculum learning strategy adaptively increases the difficulty level of selected negatives based on the learning performance of TGNNs. Additionally, these selected negatives are accompanied by annealing random negatives for robust training of TGNNs.
Adaptive Selected Negatives
Firstly, we introduce a sampling proportion $\pi_{e}$ that self-adjusts based on the model’s performance at $e$-th epoch – shrinking when performance improves and expanding otherwise – thus dynamically controlling the difficulty level of selected negatives. By starting with a large $\pi_{e}$ to include both easy and hard negatives and tightening it only if the model consistently demonstrates improvements, this approach prevents the model from being overwhelmed by overly challenging negatives. Furthermore, we utilize a shrinkage rate $\gamma_{\text{shrink}}$ to control the adjustment steps of $\pi_{e}$. A high $\gamma_{\text{shrink}}$ establishes a steep difficulty curve for model learning, while a low $\gamma_{\text{shrink}}$ sets a flatter curve. This strategy is written as
$\pi_{e+1}=\begin{cases}\pi_{e}-\gamma_{\text{shrink}},&\text{If performance improves,}\\ \pi_{e}+\gamma_{\text{shrink}},&\text{Otherwise}\end{cases}$ (1)
In addition, instead of the traditional method of selecting a set of negative samples for each positive sample, we merge and evaluate all candidate negatives at the minibatch level. As the learning difficulties vary across source nodes, this approach allows us to focus on more informative nodes that present greater difficulty while automatically eliminating nodes that are easier to learn as $\pi_{e}$ shrinks. Technically, we select the top-scored $n_{b}=\pi_{e}\times|b|\times M$ samples for each minibatch $b$ at $e$-th epoch, where $|b|$ denotes the number of source nodes in minibatch $b$, and $M$ represents the size of each source’s negative pool $V^{-}_{u}$.
$\begin{split}O^{-}_{\text{sel}}&=\text{arg top}_{n_{b}}\{\hat{s}_{uv^{-}}|\forall u\in b,v^{-}\in V^{-}_{u}\},\end{split}$ (2)
where $O^{-}_{\text{sel}}$ denotes the set of selected negatives, and $\hat{s}_{uv^{-}}$ is a scoring function between node pairs for negative selection, as defined by Eq. 13.
Annealing Random Negatives
To provide the model with additional space for exploration and adaptation, especially in early epochs, each minibatch also includes $n_{b}$ random negatives $O^{-}_{\text{rand}}$ in addition to $O^{-}_{\text{sel}}$. A weight $\delta_{e}$ is applied to balance these two types. Initially, we assign a higher weight to random negatives to enable the model to understand the basic patterns. As the model gains experience, the emphasis gradually shifts toward the selected, presumably more challenging negatives, enhancing the learning depth. Following this intuition, we define $\delta_{e}=\max\{\pi_{e},\delta_{\text{min}}\}$. $\delta_{\text{min}}$ is a hyper-parameter that controls the minimum influence random negatives should exert, ensuring that both contribute to the learning process. The loss for each minibatch is computed as
$L=L_{pos}+\delta_{e}L_{O^{-}_{\text{rand}}}+(1-\delta_{e})L_{O^{-}_{\text{sel}}}+\lambda L_{\text{con}}+\mu\lVert\Theta\rVert^{2}_{2},$ (3)
where $L_{pos}$, $L_{O^{-}_{\text{rand}}}$, and $L_{O^{-}_{\text{sel}}}$ denote the mean binary cross-entropy losses for positive samples, random negatives, and selected negatives, respectively. $L_{\text{con}}$ is the average contrastive loss as defined in Eq. 11, and $\lVert\Theta\rVert^{2}_{2}$ is a regularization term.
4.2 Negative Pool
The negative pool, which is an essential component of the curriculum learning strategy, is also constructed dynamically.
To ensure efficiency, we start the entire process by sampling a negative pool $V^{-}_{u}$ of fixed size $M$ for each positive sample $(u,v^{+},t)$. During the initial epochs, we sample $V^{-}_{u}$ as a mixture of historical neighbors and random negatives, a simple approach to prevent overcrowding with easy negatives. Random samples here serve a different purpose than before and are included to capture newly emerging preferences not disclosed by users’ historical interactions. This process can be written as
$V^{-}_{u}=V^{\text{hist}}_{u}\cup V^{\text{rand}}_{u},$ (4)
where $|V^{-}_{u}|=M$, and a proportion hyper-parameter $\gamma_{\text{hist}}$ is used to balance between $|V^{\text{hist}}_{u}|=\gamma_{\text{hist}}*M$ and $|V^{\text{rand}}_{u}|=(1-\gamma_{\text{hist}})*M$.
Once the model learns the fundamentals, we raise the difficulty level. We introduce this advanced stage when sampling proportion $\pi_{e}$ falls below a preset threshold $\tau$ for the first time. Since $\pi_{e}$ decreases only if the model is learning well, this condition ensures that the model only encounters more challenging negatives after it has adequately learned the basics. The indicator $\mathcal{C}_{\pi_{e}}$ is determined as follows:
$\mathcal{C}_{\pi_{e}}=\mathbbm{1}(\pi_{e}\leq\tau).$ (5)
In this advanced stage, we implement a hard negative cache $V^{\text{hard}}_{u}$ of size $\frac{M}{2}$ to store a subset of $V^{-}_{u}$ from the current epoch, focusing on likely false positives and those showing less learning in previous epochs. The cached negatives are sampled proportionally according to the negatives’ performance metrics, as outlined in Eq. 6.
$V^{\text{hard}}_{u}\sim\{p_{uv}-\alpha_{e}\times\text{std}(P_{uv})|\forall v\in V^{-}_{u}\},$ (6)
where $|V^{\text{hard}}_{u}|=\frac{M}{2}$, $p_{uv}$ denotes the probability score of node pair $(u,v)$ at the current epoch, $P_{uv}$ records the historical scores within the last five epochs, and std($\cdot$) is set to 00 when results are not available. $\alpha_{e}$ is a hyper-parameter that controls the importance of prediction standard deviation at the $e$-th epoch. To manage potential instability in early learning stages, we initially assign a modest weight to the standard deviation and gradually increase it. Specifically, we use $\alpha_{e}=\alpha_{\text{max}}\times\min(\frac{e}{E^{\prime}},1)$, where $E^{\prime}$ is the epoch after which the weight remains constant, and $\alpha_{\text{max}}$ is the maximum weight that standard deviations could contribute to the performance metrics.
This hard negative cache then accounts for half of the negative pool in the next epoch, while the remainder continues to be a mix of historical neighbors and random samples, maintaining the previous proportion. Therefore, the new updating strategy for each negative pool $V^{-}_{u}$ is written as
$V^{-}_{u}=V^{\text{hard}}_{u}\cup V^{\text{hist}}_{u}\cup V^{\text{rand}}_{u},$ (7)
where $\frac{|V^{\text{hist}}_{u}|}{|V^{\text{rand}}_{u}|}=\frac{\gamma_{\text{hist}}}{1-\gamma_{\text{hist}}}$ and $|V^{\text{hist}}_{u}|+|V^{\text{rand}}_{u}|=\frac{M}{2}$.
4.3 Negative Selection
The second critical component of our framework is the selection function used to choose negative samples from the negative pool $V^{-}_{u}$. This process consists of two steps: disentangle irrelevant factors of negative nodes and score negative samples with time embeddings.
Irrelevant Factor Disentanglement
To facilitate TGNNs to differentiate positive from negative samples, we disentangle nodes’ relevant and irrelevant factors, which is in line with DENS [17].
For each positive sample $(u,v^{+},t)$, we construct a positive gate. This gate leverages the embeddings of the source node $h_{u}$ and the destination node $h_{v^{+}}$ to isolate the relevant factor $h_{v^{+}}^{r}$ of the destination node $v^{+}$, illustrated as
$h_{v^{+}}^{r}=h_{v^{+}}\odot\sigma(W_{\text{src}}h_{u}+W_{\text{dst}}h_{v^{+}}+b_{p}),$ (8)
where $W_{\text{src}}$, $W_{\text{dst}}$ and $b_{p}$ are trainable parameters shared by all positive samples, $\sigma$ is the sigmoid function, and $\odot$ is the element-wise product. This factor and the embedding $h_{v^{-}}$ of the corresponding candidate negative $v^{-}\in V^{-}_{u}$ are then used as inputs to isolate the relevant factor $h_{v^{-}}^{r}$ of the negative.
$h_{v^{-}}^{r}=h_{v^{-}}\odot\sigma(W_{p}h_{v^{+}}^{r}+W_{n}h_{v^{-}}+b_{n}),$ (9)
where $W_{p}$, $W_{n}$, and $b_{n}$ are trainable parameters shared by all candidate negatives in the negative pool. Factors other than relevant factors are denoted as irrelevant factors:
$h_{v^{+}}^{ir}=h_{v^{+}}\ominus h_{v^{+}}^{r},\quad h_{v^{-}}^{ir}=h_{v^{-}}\ominus h_{v^{-}}^{r},$ (10)
where $h_{v^{+}}^{ir}$ and $h_{v^{-}}^{ir}$ denote irrelevant factors of the destination node $v^{+}$ and the candidate negative $v^{-}$, respectively, and $\ominus$ is element-wise subtraction.
To supervise the training of positive and negative gates, we introduce a four-component contrastive loss $L_{\text{con}}$. These components dictate that relevant factors mainly drive interactions in positive samples and the absence of interactions in negative samples, while irrelevant factors do not and may even favor negative samples. This approach is consistent with the procedure used in DENS [17], written as
$L_{\text{con}}=-(s_{r,ir}^{+,+}+s_{r,r}^{+,-}+s_{ir,r}^{-,-}+s_{ir,ir}^{-,+}),$ (11)
where $s_{g1,g2}^{sgn1,sgn2}=MLP(h_{u},h_{v^{sgn1}}^{g1})-MLP(h_{u},h_{v^{sgn2}}^{g2})$ is the scoring function between different factors, with $MLP(\cdot,\cdot)$ computing the probability score from the given embeddings, and a minus sign is added to perform maximization.
Temporal Scoring Function
As our goal is for the model to predict neighbors based on relevant factors, the most useful negatives for training are those that differ from the positives only in their relevant factors. In practice, this involves selecting negative samples from the negative pool $V^{-}_{u}$ by maximizing the differences in relevant factors and minimizing the differences in irrelevant factors. This approach is written as
$s_{uv^{-}}$ $=$ $\beta_{e}\times s_{r,r}^{+,-}-(1-\beta_{e})\times s_{ir,ir}^{+,-},$ (12)
where $s_{uv^{-}}$ denotes the scoring function and $\beta_{e}$ is a hyper-parameter that balances the contributions of relevant and irrelevant factors at epoch $e$. Intuitively, we want the model to first learn which factors are important and then understand how these factors drive interactions. This entails initially focusing on aligning irrelevant factors and later shifting to differentiating relevant factors. Specifically, we set $\beta_{e}=\min(\frac{e}{E^{\prime\prime}},1)$, where $E^{\prime\prime}$ marks the epoch from which the score is solely determined by relevant factors.
Additionally, to capture the unique aspects of temporal networks, we incorporate three Time2Vec models to represent complex temporal patterns, including recurrent and periodic behaviors. Therefore, the final temporal scoring function used for selecting negative samples is written as
$\hat{s}_{uv^{-}}=t_{u,v^{+}}\ast(t_{u,v^{-}}\odot t_{v^{-}})\odot s_{uv^{-}},$ (13)
where $t_{u,v^{+}}$, $t_{u,v^{-}}$, and $t_{v^{-}}$ are Time2Vec embeddings for the timestamps of the positive sample, the interaction between $u$ and the candidate negative $v^{-}\in V^{-}_{u}$, and the negative’s most recent occurrence, respectively. Default inputs of 0 are provided into Time2Vec for cases where the negative is not a historical neighbor or if it never occurs. By incorporating the product of these three embeddings, we seek to prioritize historical neighbors and recently popular nodes. The rationale is that these negatives possess attributes that were favored by the source nodes not long ago, and they might trick the model. So, by assigning more weight to these negatives – particularly those that differ from positives only in relevant factors – we can enhance the precision of the model’s decision boundary and reduce the occurrence of false negatives.
5 Experiment
5.1 Datasets and Evaluation
Dataset
We evaluate the performance of our CurNM method on $12$ standardized datasets obtained from DyGLib [35] and TGL [40]. These datasets encompass a broad spectrum of domains, sizes, durations, and network features, and are intended to represent a diverse array of datasets. According to Table 1, which summarizes the statistics, social and interaction networks are typically sparse, while social and economic networks often exhibit high recency and repetition. These characteristics are expected due to the limited number of relationships individuals can realistically maintain and the significant impact of recent events and cyclic patterns on public attention and market dynamics. Other features, such as average time proximity, present a more diverse result across different domains. We divide these datasets into training, validation, and test sets using a chronological split with ratios of $70:15:15$ to prevent temporal data leakage during model training.
Table 1: Statistical Overview of the Datasets. Average Proximity is the average step difference between nodes and their historical neighbors, Recency is the percentage of current interactions that replicate the last, Repetition is the proportion of repeated edges, $d_{\text{max}}$ represents the maximum temporal node degree.
Dataset Domain Bipartite $|V|$ $|E|$ Average Proximity Recency Repetition Density $d_{\text{max}}$ Time Span
uci Social No 1900 59835 58 34% 77% 5.63E-07 1.55E+03 194 days
USLegis Politics No 226 60396 103 40% 72% 1.98E-01 1.20E+03 11 congresses
CanParl Politics No 735 74478 34 25% 39% 1.97E-02 2.38E+03 13 years
enron Social No 185 125235 1004 45% 98% 3.23E-04 2.15E+04 4 years
WIKI Social Yes 9228 157474 62 79% 88% 2.42E-08 1.94E+03 1 month
MOOC Interaction No 7047 411749 30 20% 57% 4.80E-08 1.95E+04 1 month
UNtrade Economics No 256 507497 890 83% 96% 4.84E-01 1.22E+04 31 years
REDDIT Social Yes 10985 672447 67 61% 88% 1.67E-08 5.87E+04 1 month
myket Interaction Yes 17989 694121 26 5% 21% 6.18E-09 1.53E+04 6.5 months
UNvote Politics No 202 1035742 1570 92% 98% 7.09E-01 1.59E+04 71 years
LASTFM Interaction Yes 1980 1293103 1713 10% 88% 5.14E-07 5.18E+04 1 month
Flights Transport No 13170 1927145 822 40% 84% 1.82E-04 3.47E+04 1 month
Evaluation Metrics
We use average precision (AP) as our primary evaluation metric. Specifically, we compare AP for each dataset across all combinations of TGNNs and sampling methods, since typically only the optimal setup is used in practice. Additionally, given that we conduct experiments across 12 datasets, we evaluate the overall performance of each sampling method using average rank.
These two metrics are assessed across two scenarios: transductive and inductive negative evaluation. In the transductive setting, our test includes a mix of half-random and half-historical (TRANS) negative samples. Given that our method is specifically tailored to enhance performance against hard negatives, we anticipate improved predictive accuracy for historical neighbors [35]. However, it is crucial to preserve the model’s capability to accurately predict random negatives. In this context, TRANS provides a robust benchmark for assessing our method’s ability to balance these two objectives. Conversely, the inductive test involves only nodes not seen during training and thus provides a valuable indicator of our method’s generalizability to new data. This evaluation draws on code from DyGLib [35].
5.2 Experiment Setting
Baselines
We compare our method against two commonly adopted negative sampling strategies – random sampling and recent sampling [24, 33] – along with ENS [7], the only existing method known to us tailored for temporal networks. Random sampling selects negative samples based on a uniform distribution. Recent sampling selects the latest historical neighbors as negatives, offering a straightforward approach to addressing issues of positive sparsity and positive shift. Finally, ENS trains models using a predetermined proportion of random negatives and historically appearing nodes while ensuring that the average temporal proximity reaches a specified level.
We conduct our experiments on JODIE [16], TGN [24], and DySAT [25], three well-established models that exhibit superior performance on our datasets with random sampling.
Implementation Details
For a fair comparison, we follow default implementations in TGL [40]. In all experiments, the Adam optimizer is used with a learning rate of $0.0001$ and a mini-batch size of $200$. We perform each experiment for up to $400$ epochs and repeat it with $5$ different random seeds. All methods are executed with a 48-core CPU, 128GB of memory, and a 24G P5000 GPU.
The contrastive loss coefficient $\lambda$ and the $L_{2}$ regularization coefficient $\mu$ are set to ensure a balanced treatment across different loss components. The remaining hyperparameters are carefully tuned through grid search. The size of the negative pool $M$ is tested within $\{4,8,12,16\}$; the proportion of historical neighbors $\gamma_{\text{hist}}$ within $\{0.3,0.4,0.5,0.6,0.7\}$; and the hard negative cache’s initiation threshold $\tau$ within $\{0.3,0.5,0.7\}$. The stopping epoch $E^{\prime}$ and the maximum weight $\alpha_{\text{max}}$ for $\alpha_{e}$ are explored over $\{\frac{E_{0}}{10},\frac{E_{0}}{2},E_{0}\}$ and $\{0,0.006,0.012,0.024\}$ respectively, where $E_{0}$ denotes the maximum number of epochs. Similarly, $E^{\prime\prime}$ for $\beta_{e}$ and $\delta_{\text{min}}$ for $\delta_{e}$ are searched within $\{\frac{E_{0}}{10},\frac{E_{0}}{2},E_{0}\}$ and $\{0.3,0.5,0.7\}$. All chosen values are included in the code. We will elaborate on tests involving $\alpha_{e}$, $\beta_{e}$, $\pi_{e}$, and $\delta_{e}$ in a later section.
5.3 Experimental Results
As shown in Table 2, our method consistently achieves the best ranks in both historical and inductive settings, underscoring its efficacy.
Table 2: Average Precision of link prediction across 12 datasets and three TGNNs. bold denotes the best performance achieved for each dataset, - indicates failed results, Avg Rank refers to the average ranks of each sampling method.
Dataset Model Transductive Inductive
Random Recent ENS CurNM Random Recent ENS CurNM
uci DySAT 0.763 0.787 0.763 0.764 0.726 0.727 0.737 0.741
JODIE 0.839 0.564 0.855 0.848 0.636 0.542 0.648 0.640
TGN 0.824 0.651 0.849 0.832 0.677 0.662 0.700 0.695
USLegis DySAT 0.505 0.602 0.564 0.539 0.480 0.584 0.510 0.517
JODIE 0.492 0.527 0.501 0.515 0.597 0.596 0.618 0.620
TGN 0.808 0.797 0.833 0.868 0.763 0.730 0.789 0.809
CanParl DySAT 0.586 0.539 0.597 0.577 0.672 0.630 0.674 0.664
JODIE 0.572 0.569 0.570 0.652 0.533 0.539 0.530 0.570
TGN 0.747 0.705 0.713 0.791 0.626 0.624 0.604 0.650
enron DySAT 0.744 0.727 0.742 0.742 0.704 0.698 0.703 0.706
JODIE 0.800 0.752 0.801 0.826 0.724 0.693 0.716 0.739
TGN 0.771 0.842 0.824 0.848 0.730 0.811 0.769 0.787
WIKI DySAT 0.836 0.844 0.843 0.887 0.821 0.804 0.823 0.834
JODIE 0.791 0.734 0.786 0.818 0.701 0.701 0.712 0.728
TGN 0.906 0.883 0.908 0.917 0.856 0.837 0.860 0.868
MOOC DySAT 0.699 0.611 0.720 0.710 0.684 0.592 0.686 0.681
JODIE 0.670 0.564 0.679 0.679 0.580 0.574 0.576 0.577
TGN 0.764 0.621 0.774 0.767 0.688 0.653 0.663 0.690
UNtrade DySAT 0.527 0.534 0.528 0.527 0.538 0.556 0.539 0.542
JODIE 0.609 0.655 0.603 0.613 0.609 0.680 0.610 0.619
TGN 0.644 0.660 0.647 0.665 0.661 0.665 0.660 0.681
REDDIT DySAT 0.835 0.813 0.842 0.840 0.823 0.814 0.829 0.819
JODIE 0.823 0.781 0.835 0.829 0.787 0.777 0.794 0.797
TGN 0.878 0.848 0.890 0.878 0.874 0.853 0.871 0.877
myket DySAT 0.668 0.738 0.665 0.667 0.596 0.689 0.592 0.596
JODIE 0.674 0.664 0.685 0.688 0.530 0.533 0.545 0.537
TGN 0.694 0.679 0.692 0.702 0.582 0.539 0.581 0.588
UNvote DySAT 0.510 0.502 0.508 0.516 0.511 0.499 0.509 0.519
JODIE 0.684 0.639 0.696 0.714 0.694 0.635 0.698 0.703
TGN 0.681 0.601 0.693 0.709 0.650 0.594 0.660 0.673
LASTFM DySAT 0.701 0.640 0.704 0.754 0.709 0.622 0.712 0.738
JODIE 0.536 0.549 0.551 0.558 0.549 0.526 0.558 0.571
TGN 0.742 0.603 0.740 0.700 0.706 0.596 0.701 0.676
Flights DySAT 0.737 0.699 - 0.511 0.710 0.689 - 0.521
JODIE 0.736 0.484 - 0.772 0.656 0.605 - 0.689
TGN 0.771 0.573 - 0.807 0.696 0.665 - 0.747
Avg Rank 2.75 3.19 2.18 1.72 2.64 3.28 2.39 1.56
Transductive Setting
In the transductive setting, our method delivers the best results across 8 datasets, with its overall average rank nearly half a point better than the second-best method, ENS, and more than a point ahead of random and recent samplings. This suggests that our method outperforms all existing methods in predicting historical recurrence while maintaining strong performance in predicting random samples.
At the dataset level, our method excels in relatively dense networks with frequent recurrence of recent historical interactions (low Average Time Proximity), such as USLegis, CanParl, and Flights, where it surpasses the best existing methods by around 5%. This demonstrates that our method, by utilizing temporal-aware disentangled factors for selection, can more accurately distinguish between enduring and transient relationships over time. In networks with high repetition and recency effects – typically favorable for recent sampling – such as enron, UNtrade, and UNvote, our method still outperforms recent sampling. Furthermore, it achieves an average improvement of nearly 5% over the best random sampling and nearly 3% over ENS, showcasing its effectiveness in capturing periodic and recurrent patterns. These results confirm our method’s capability in addressing positive shift.
In very sparse networks, our method also outperforms random sampling across all datasets and, on average, surpasses ENS and recent sampling. Even where it does not lead, it remains competitively close to the top performer. The only exception is myket, whose sparsity, recency, and repetition make it akin to a static network, rendering ENS and CurNM – methods designed for temporal networks – less effective. Despite this, the overall performance indicates that our method successfully circumvents the steep gradient curve prevalent in hard negative sampling, effectively addressing positive sparsity even in less favorable scenarios.
Inductive Setting
The inductive setting allows us to compare our method’s generalizability against existing methods by evaluating their performances on unseen nodes. When comparing inductive to transductive settings, the average ranks of ENS and recent sampling worsen, suggesting that their heuristic strategies might not be generalizable and could lead to overfitting on seen nodes. On the contrary, the average rank of random sampling improves, indicating its robustness despite relatively weaker performance. Finally, our CurNM method consistently achieves the best across 9 datasets and delivers a higher average rank in the inductive setting compared to the transductive setting. This confirms that our method not only performs well but is also robust and highly generalizable.
At the dataset level, our method consistently outperforms or matches the existing methods in almost all datasets, with insignificant declines in only two cases (again, with myket being an outlier).
5.4 Complexity Analysis
Theoretically, CurNM adds a negative mining process with a time complexity of $O(|E|\times n\times d)$ to TGNNs, which originally have a complexity of $O(L\times|E|\times n\times d)$. Here, $L$ denotes the number of layers, $|E|$ is the number of edges, and $n$ represents the number of positive and negative samples per node, with nonlinear operations simplified to $O(d)$.
Empirically, CurNM maintains high efficiency by limiting $n$ with a relatively small negative pool, sized at $M=8$. Table 3 presents a runtime comparison between CurNM and baseline methods on the TGN model, which demonstrates the best overall performance among the three TGNNs evaluated. According to Table 3, CurNM adds minimal computational load compared to basic random and recent sampling, consistently requiring about twice the time regardless of dataset size. Moreover, it is significantly more efficient than the state-of-the-art ENS on large datasets, requiring approximately one-seventh of the time on Flights, the largest dataset tested.
Table 3: Comparison of average runtimes per epoch between CurNM and baseline methods on the TGN model. $Ratio_{Rand}$, $Ratio_{Rec}$, and $Ratio_{ENS}$ represent the ratio of CurNM runtime to Random, Recent, and ENS runtimes, respectively.
Dataset Random $Ratio_{Rand}$ Recent $Ratio_{Rec}$ ENS $Ratio_{ENS}$ CurNM
uci 8.0 1.86 8.6 1.73 10.8 1.38 14.9
USLegis 8.8 1.86 8.6 1.91 12.1 1.36 16.4
CanParl 11.3 1.47 11.4 1.46 17.1 0.97 16.6
enron 16.4 1.7 17.0 1.64 17.7 1.58 27.9
WIKI 22.7 1.45 23.4 1.41 28.5 1.15 32.9
MOOC 39.9 2.09 43.4 1.92 152.7 0.55 83.3
UNtrade 65.6 1.83 71.8 1.67 154.6 0.78 120.1
REDDIT 84.9 1.76 116.7 1.28 290.8 0.51 149.0
myket 79.3 2.18 89.2 1.94 1133.9 0.15 172.8
UNvote 116.0 1.91 138.4 1.6 422.3 0.52 221.0
LASTFM 161.2 2.29 156.2 2.36 1019.0 0.36 368.8
Flights 221.8 2.1 231.7 2.01 3222.5 0.14 466.6
5.5 Ablation Study
The ablation study detailed in Figure 2 examines the effects of adaptive $\pi_{e}$, annealing random negatives, hard negative cache, factor disentanglement, and temporal-aware embeddings in our method on the TGN model. Due to space constraints, we present our study using six datasets here. In both transductive and inductive settings, the complete CurNM model delivers the most robust performance across the datasets. The most significant decline in performance is observed when random negatives are removed, validating our claim that a steep difficulty curve can confuse the model. In datasets like USLegis, CanParl, and enron, where our model achieves strong results, the removal of temporal-aware embeddings and adaptive $\pi_{e}$ also leads to substantial performance drops. This underscores their critical role in addressing the networks’ positive sparsity and positive shift. Although the impact of removing disentanglement and hard negative cache is less pronounced, their absence can still result in up to a 3% performance decline, suggesting their importance in enhancing the granularity of learning within the model.
Figure 2: Ablation studies on different components.
5.6 Hyper-Parameter Sensitivity
Impact of Strategies for $\alpha_{e},\beta_{e}$, and $\delta_{e}$
We investigate our assumptions about the structures of $\alpha_{e},\beta_{e}$, and $\delta_{e}$ by testing an increasing strategy (e.g., $\alpha_{e}=\min(\frac{e}{E^{\prime}},1)$), a decreasing strategy (e.g., $\alpha_{e}=\max(0,1-\frac{e}{E^{\prime}})$), and a constant strategy (e.g., $\alpha_{e}=0.5$). Our findings in Figure 3 confirm that, as expected, our current approach yields the most robust results. This underscores the effectiveness of our dynamic strategies, which progressively guide the model to focus on relevant factors of informative nodes, thereby achieving a smoother and more advantageous learning curve.
Impact of Strategies for $\pi_{e}$ and $\gamma_{\text{shrink}}$
Under the framework where $\pi_{e}$ should shrink as learning progresses, we explore two approaches: a linear strategy and an adaptive strategy. The linear strategy consistently reduces the $\pi_{e}$ by a set amount each epoch, whereas the adaptive strategy only does so when there’s an improvement in the selected performance metric. For our baseline, both methods start with $\pi_{e}=100\%$, with a shrinkage rate $\gamma_{\text{shrink}}=3\%$, and the $\pi_{e}$ can decrease to a minimum of $10\%$. The choice of $\gamma_{\text{shrink}}=3\%$ is based on that models generally stabilize after approximately $30$ epochs, and having a larger $\pi_{e}$ before this phase to capture more information would be beneficial. The adaptive strategy outperforms the linear one. We then experiment with a slower ($\gamma_{\text{shrink}}=2\%$) and a faster shrinkage rate ($\gamma_{\text{shrink}}=5\%$). Figure 3 shows that $\gamma_{\text{shrink}}=3\%$ yields the best results.
Figure 3: Parameter sensitivity studies on $\alpha_{e},\beta_{e},\delta_{e},\text{ and }\pi_{e}$.
6 Conclusion
In this paper, we identify two unique challenges of negative sampling in temporal networks – positive sparsity and positive shift. Our proposed CurNM method employs a dynamically updated negative pool and a curriculum learning strategy to adaptively adjust the difficulty level of negative samples during training, thus addressing the common challenge of positive sparsity. Furthermore, we implement a temporal-aware negative selection function to effectively capture positive shift. Our method’s efficacy is demonstrated through extensive experiments against 3 algorithms and 12 datasets. Although CurNM demonstrates superior performance, it requires tuning several hyper-parameters to optimize its performance. In the future, we aim to simplify these designs to reduce the need for extensive parameter tuning.
Declaration of interests
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Acknowledgements
This work was supported by the Joint Funds of the Zhejiang Provincial Natural Science Foundation of China [grant number LHZSD24F020001]; Zhejiang Province ”JianBingLingYan+X” Research and Development Plan [grant number 2024C01114]; Ningbo Natural Science Foundation [grant number 2023J281]; and Zhejiang Province High-Level Talents Special Support Program ”Leading Talent of Technological Innovation of Ten-Thousands Talents Program” [grant number 2022R52046].